### v1

@@ 4.3 Overall Computational Analysis
In this section, we illustrate how all of the above results can be put together to achieve low computational cost for Kaczmarz++. For clarity, we make a few simplifying assumptions: we let ${\mathbf{A}}$ have full row rank (although similar claims can be recovered for any consistent linear system), and we allow optimal selection of the algorithmic hyper-parameters. Note that all of our intermediate results show that the algorithm is robust to different choices of parameters $\lambda$, $\rho$, $\eta$, and $B$, and in the following section, we discuss how to select these parameters in practice.
@@
Given matrix ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$ with rank $m\leq n$, and ${\mathbf{b}}\in{\mathds{R}}^{n}$, let ${\mathbf{x}}^{*}$ be the minimum-norm solution of ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$. For $\delta\in(0,1)$ and $C\log(m/\delta)\leq k<m$, Kaczmarz++ (Alg. 1) with block size $s=\lceil Ck\log(m\bar{\kappa}_{k})\rceil$, $\lambda=\frac{1}{m}\sum_{i>k}\sigma_{i}^{2}({\mathbf{A}})$, $\rho=\frac{k}{12m\bar{\kappa}_{k}}$, $\eta=\frac{k}{16m}$, ${\mathbf{x}}_{0}=\mathbf{0}_{n}$, and $B=\lceil C\frac{m}{k}\log(m/\delta)\rceil$, after $t=O(\frac{m}{k}\bar{\kappa}_{k}\log 1/\epsilon\delta)$ iterations, with probability $1-\delta$ satisfies
$\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|\leq\epsilon\|{\mathbf{x}}^{*}\|\quad\text{using}\quad\tilde{O}\left(mk^{2}+mn\bar{\kappa}_{k}\log 1/\epsilon\delta\right)\quad\text{operations.}$.
Remark 4.6.
Similar guarantee (up to logarithmic factors) can be recovered for Kaczmarz++ running on a consistent over-determined linear system (i.e., $m\geq n$).
@@
Corollary 4.4 implies that it takes $t=O(\frac{m}{k}\bar{\kappa}_{k}\log 1/\epsilon)$ iterations to converge $\epsilon$-close in expectation. We convert this to a guarantee that holds with $1-\delta$ probability by replacing $\epsilon$ with $\epsilon\delta$ and then applying Markov’s inequality.
@@
    Then, computing all of the Cholesky factors takes $O(B(ns\log n+s^{3}\log s))=O(mn\log^{3}(m\bar{\kappa}_{k}/\delta)+mk^{2}\log k\log^{4}(m\bar{\kappa}_{k}))$.
@@
    Finally, each iteration of the algorithm requires solving (4.1) with LSQR so that $\|\tilde{\mathbf{w}}_{t}-{\mathbf{w}}_{t}\|\leq\epsilon\|{\mathbf{w}}_{t}\|$ for $\epsilon\leq\frac{k}{Cm\bar{\kappa}_{k}^{2}}$, which for a given block $S$ takes $O(ns\log(\kappa_{S}/\epsilon))$, where $\kappa_{S}=\sqrt{\lambda}/\sigma_{\min}({\mathbf{A}}_{S})$. Note that since ${\mathbf{A}}$ has full row rank, $\sigma_{\min}({\mathbf{A}}_{S})\geq\sigma_{\min}({\mathbf{A}})$, so thanks to our choice of $\lambda$, we have $\kappa_{S}\leq\bar{\kappa}_{k}$ for any $S$. Thus, the cost of each iteration is $O(nk\log^{2}(n\bar{\kappa}_{k}))$. Multiplying by $t$, we get $O(mn\bar{\kappa}_{k}\log^{2}(n\bar{\kappa}_{k})\log(1/\epsilon\delta))$ operations.
@@ 5.1 Coordinate Descent
When the matrix ${\mathbf{A}}$ is PSD, then Kaczmarz++ admits a specialized formulation as a block coordinate descent method, similarly as can be done for the classical randomized Kaczmarz algorithm, see e.g. [HNR17, Pet15]. To see this, let $\tilde{\mathbf{x}}=\mathcal{A}({\mathbf{A}},{\mathbf{b}},{\mathbf{x}}_{0})$ denote the output of Algorithm 1 for solving a general linear system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ (for now, skipping the RHT step), initialized with ${\mathbf{x}}_{0}$. Now, if we are given an $n\times n$ PSD matrix, ${\mathbf{A}}\in\mathcal{S}_{n}^{+}$, then we can write ${\mathbf{A}}={\mathbf{\Phi}}{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}$ for some ${\mathbf{\Phi}}$ (for example, the square root of ${\mathbf{A}}$), and so we can rewrite ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ as:
${\mathbf{\Phi}}{\mathbf{z}}={\mathbf{b}}\quad\text{for}\quad{\mathbf{z}}={\mathbf{\Phi}}^{\scriptscriptstyle{\top}}{\mathbf{x}}.$
Thus, instead of solving the system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ directly, we consider implicitly applying our block Kaczmarz algorithm to ${\mathbf{\Phi}}{\mathbf{z}}={\mathbf{b}}$, i.e., computing ${\mathbf{\Phi}}^{\scriptscriptstyle{\top}}\tilde{\mathbf{x}}=\mathcal{A}({\mathbf{\Phi}},{\mathbf{b}},{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}{\mathbf{x}}_{0})$. For this procedure, letting ${\mathbf{z}}_{t}={\mathbf{\Phi}}^{\scriptscriptstyle{\top}}{\mathbf{x}}_{t}$ be the implicit Kaczmarz++ iterates, we get:
${\mathbf{w}}_{t}={\mathbf{\Phi}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{\Phi}}_{S}{\mathbf{\Phi}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}({\mathbf{\Phi}}_{S}{\mathbf{z}}_{t}-{\mathbf{b}}_{S})={\mathbf{\Phi}}^{\scriptscriptstyle{\top}}{\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S,S}+{\lambda}{\mathbf{I}})^{-1}({\mathbf{A}}_{S}{\mathbf{x}}_{t}-{\mathbf{b}}_{S}).$ (5.1)
This leads to the following coordinate descent update (line 11 in Algorithm 3):
${\mathbf{w}}_{t}^{\mathrm{cd}}$ $={\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S,S}+{\lambda}{\mathbf{I}})^{-1}({\mathbf{A}}_{S}{\mathbf{x}}_{t}-{\mathbf{b}}_{S}).$
@@
Thus, the convergence rate of CD++ is determined by the convergence of the implicit Kaczmarz++ algorithm, since $\|{\mathbf{z}}_{t}-{\mathbf{z}}^{*}\|=\|{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}({\mathbf{x}}_{t}-{\mathbf{x}}^{*})\|=\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|_{{\mathbf{A}}}$. Crucially, this convergence is now governed by $\bar{\kappa}_{k}({\mathbf{\Phi}})\leq\bar{\kappa}_{k}({\mathbf{A}})$, hence the improvement over simply running Kaczmarz++ on ${\mathbf{A}}$.
@@ 5.2 Simplified Block Memoization
This block selection strategy implies that the first $B=\frac{n}{s}\log n$ blocks will be sampled uniformly at random from all size $s$ index sets and added to the block list ${\mathcal{B}}$. After that, in $T$ iterations the scheme will collect on average an additional $\sum_{t=B}^{T}\frac{B}{t}\approx B\log(\frac{T}{B})$ blocks. Since the factor $\log(\frac{T}{B})$ grows as the iterations progress, this implies that we are guaranteed to reach the number of blocks that is needed by our theory to imply convergence. However, since the factor grows slowly, we will not have to do too much unnecessary Cholesky factorizations before converging to a desired accuracy. Note that each Cholesky factor takes $O(s^{2})$ memory, so if we store $O(\frac{n}{s}\log n)$ factors throughout the convergence, then this uses only $O(ns\log n)$ additional memory.

### v2

@@ 4.3 Overall Computational Analysis
In this section, we illustrate how all of the above results can be put together to achieve low computational cost for Kaczmarz++, while ensuring fast convergence towards the optimum.
We start by combining Theorem 4.2 with the analysis of $\mu$ and $\nu$ in Theorems 3.1 and 3.8, as well as the acceleration analysis from Theorem 2.1, in order to recover the convergence guarantee for Kaczmarz++ (Algorithm 1).
Corollary 4.4.
Given ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$ with $\mathrm{rank}({\mathbf{A}})=r$ such that ${\mathbf{A}}{\mathbf{x}}^{*}={\mathbf{b}}$, let $\delta\in(0,r/m\bar{\kappa}_{k}^{2})$ and $C\log(m/\delta)\leq k<r$. There are $\lambda,\rho,\eta$ such that if Algorithm 1 solves the regularized projection step (line 8) via Lemma 4.1 with $\epsilon\leq\frac{\rho^{2}}{8\eta}$ and samples $B\geq C\frac{r}{k}\log(mn\bar{\kappa}_{k})\log(n/\delta)$ blocks, then conditioned on a $1-\delta$ probability event depending only on the RHT ${\mathbf{Q}}$ and block set ${\mathcal{B}}$, we have:
${\mathds{E}}\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}\leq 8\left(1-\frac{k}{24\tilde{r}}\bar{\kappa}_{k}\right)^{t}\cdot\|{\mathbf{x}}_{0}-{\mathbf{x}}^{*}\|^{2}\quad\text{where}\quad\tilde{r}=\min\Big{\{}m,3r\sqrt{\log(mn\bar{\kappa}_{k})}\Big{\}}.$
Proof.
Choose $\lambda=\frac{1}{m}\sum_{i>k}\sigma_{i}^{2}({\mathbf{A}})$ and let ${\mathcal{U}}({\mathcal{B}})$ denote the uniform distribution over the block sets. Also, let $\bar{\mathbf{A}}={\mathbf{Q}}{\mathbf{A}}$ denote the input matrix after RHT preprocessing. Conditioned on the $1-\delta$ probability event defined in Corollary 4.3, by applying Theorem 3.1 with ${\mathcal{D}}={\mathcal{U}}({\mathcal{B}})$ we obtain the following:
$\mu\big{(}\bar{\mathbf{A}},{\mathcal{U}}({\mathcal{B}}),\lambda\big{)}\geq\frac{k}{4r\bar{\kappa}_{k}^{2}}.$
Moreover, according to Corollary 4.3, for any $i\in[B]$, conditioned on the same event we have ${\mathbf{A}}_{S_{i}}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S_{i}}\preceq\min\{1,\frac{9r\log(mn\bar{\kappa}_{k})}{m}\}{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}$ for all $i\in[B]$. Thus, by Theorem 3.8,
$\nu\big{(}\bar{\mathbf{A}},{\mathcal{U}}({\mathcal{B}}),\lambda\big{)}\leq\min\left\{\frac{8m}{k},\frac{80r}{k}\log(mn\bar{\kappa}_{k})\right\}.$
Combining the above bounds on $\mu$ and $\nu$, we obtain the convergence rate as
$\bar{\rho}\big{(}\bar{\mathbf{A}},{\mathcal{U}}({\mathcal{B}}),\lambda\big{)}=\sqrt{\frac{\mu\big{(}\bar{\mathbf{A}},{\mathcal{U}}({\mathcal{B}}),\lambda\big{)}}{\nu\big{(}\bar{\mathbf{A}},{\mathcal{U}}({\mathcal{B}}),\lambda\big{)}}}\geq\frac{k}{6\tilde{r}\bar{\kappa}_{k}}.$
We conclude the proof via Theorem 2.1 by choosing $\rho=\bar{\rho}/2$ and $\eta=\frac{1}{2\nu}$. ∎
We are now ready to provide the overall computational analysis of our algorithm. While we allow optimal selection of the algorithmic hyper-parameters in this statement, note that all of our intermediate results show that the algorithm is robust to different choices of parameters $\lambda$, $\rho$, $\eta$, and $B$, and in the following section, we discuss how to efficiently select these parameters during runtime.
@@
Given ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$ with $\mathrm{rank}({\mathbf{A}})=r$ and ${\mathbf{b}}\in{\mathds{R}}^{m}$, let ${\mathbf{x}}^{*}$ be the minimum-norm solution of ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$. For $\delta\in(0,1)$ and $C\log(m/\delta)\leq k<r$, Kaczmarz++ (Alg. 1 via Corr. 4.4) with $s=\lceil Ck\log(m\bar{\kappa}_{k})\rceil$, $\lambda=\frac{1}{m}\sum_{i>k}\sigma_{i}^{2}({\mathbf{A}})$, $\rho=\frac{k}{24r\bar{\kappa}_{k}\log^{1/2}(mn\bar{\kappa}_{k})},B=\lceil\frac{Cr}{k}\log(mn\bar{\kappa}_{k})\log(2n/\delta)\rceil$, and $\eta=\frac{k}{160r\log(mn\bar{\kappa}_{k})}$, ${\mathbf{x}}_{0}=\mathbf{0}_{n}$, after $t=\tilde{O}(\frac{r}{k}\bar{\kappa}_{k}\log 1/\epsilon\delta)$ iterations, with probability $1-\delta$ has
$\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|\leq\epsilon\|{\mathbf{x}}^{*}\|\quad\text{using}\quad\tilde{O}\left(mn+rk^{2}+nr\bar{\kappa}_{k}\log 1/\epsilon\delta\right)\quad\text{operations.}$.
@@
Corollary 4.4 implies that it takes $t=\tilde{O}(\frac{r}{k}\bar{\kappa}_{k}\sqrt{\log(mn\bar{\kappa}_{k})}\log 1/\epsilon)$ iterations to converge $\epsilon$-close in expectation. We convert this to a guarantee that holds with $1-\delta$ probability by replacing $\epsilon$ with $\epsilon\delta$ and then applying Markov’s inequality.
@@
    Then, computing all of the Cholesky factors takes $O(B(ns\log n+s^{3}\log s))=\tilde{O}(nr+rk^{2})$.
@@
    Finally, each iteration of the algorithm requires solving (4.1) with LSQR so that $\|\tilde{\mathbf{w}}_{t}-{\mathbf{w}}_{t}\|\leq\epsilon^{\prime}\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|$ for $\epsilon^{\prime}\leq\frac{k}{Cr\bar{\kappa}_{k}^{2}}$, which for a given block $S$ takes $O(ns\log(1/\epsilon^{\prime}))$ operations. Thus, the cost of each iteration is $\tilde{O}(nk)$. Multiplying by $t$, we get $\tilde{O}(nr\bar{\kappa}_{k}\log 1/\epsilon\delta)$ operations.
@@ 5.1 Coordinate Descent
When the matrix ${\mathbf{A}}$ is PSD, then Kaczmarz++ admits a specialized formulation as a block coordinate descent method, similarly as can be done for the classical randomized Kaczmarz algorithm, see e.g. [HNR17, Pet15]. By applying our Kaczmarz++ result to ${\mathbf{A}}^{1/2}$ we have the following convergence guarantee.
Theorem 5.1.
Suppose a PSD matrix ${\mathbf{A}}\in\mathcal{S}_{n}^{+}$ and ${\mathbf{b}}\in{\mathds{R}}^{n}$ are transformed by RHT so that $\bar{{\mathbf{A}}}={\mathbf{Q}}{\mathbf{A}}{\mathbf{Q}}^{\scriptscriptstyle{\top}}$ and $\bar{\mathbf{b}}={\mathbf{Q}}{\mathbf{b}}$, and let ${\mathbf{x}}^{*}$ be the minimum-norm solution to ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$. Given $\delta\in(0,1)$ and $C\log(n/\delta)\leq k<\mathrm{rank}({\mathbf{A}})$, let ${\mathcal{B}}$ consist of $B$ uniformly random sets from (${[n]\choose s}$ and ${\mathcal{U}}({\mathcal{B}})$ be the uniform distribution over ${\mathcal{B}}$. If we run the following update (CD++, see lines 10-13 of Algorithm 3):
$\begin{cases}{\mathbf{w}}_{t}\leftarrow{\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}(\bar{\mathbf{A}}_{S,S}+\lambda{\mathbf{I}})^{-1}(\bar{\mathbf{A}}_{S}{\mathbf{x}}_{t}-\bar{\mathbf{b}}_{S})\text{ for }S\sim{\mathcal{U}}({\mathcal{B}})\\ {\mathbf{m}}_{t+1}\leftarrow\frac{1-\rho}{1+\rho}\big{(}{\mathbf{m}}_{t}-{\mathbf{w}}_{t}\big{)}\\ {\mathbf{x}}_{t+1}\leftarrow{\mathbf{x}}_{t}-{\mathbf{w}}_{t}+\eta{\mathbf{m}}_{t+1}\end{cases}$ (5.1)
initialized with ${\mathbf{x}}_{0}\in{\mathds{R}}^{n}$, block size $s=\lceil Ck\log(n\bar{\kappa}_{k}({\mathbf{A}}^{1/2}))\rceil$, $\lambda=\frac{1}{n}\sum_{i>k}\lambda_{i}({\mathbf{A}})$, and the number of blocks $B=\lceil C\frac{n}{k}\log(n/\delta)\rceil$, then we have
${\mathds{E}}\left\|{\mathbf{Q}}^{\scriptscriptstyle{\top}}{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\right\|_{{\mathbf{A}}}^{2}\leq 8\Big{(}1-\frac{k}{24n\bar{\kappa}_{k}({\mathbf{A}}^{1/2})}\Big{)}^{t}\cdot\|{\mathbf{Q}}^{\scriptscriptstyle{\top}}{\mathbf{x}}_{0}-{\mathbf{x}}^{*}\|_{{\mathbf{A}}}^{2}.$
Proof.
We start by describing the reduction from CD++ to Kaczmarz++. For the sake of notation, suppose that ${\mathbf{x}}_{t}=\mathcal{A}_{t}({\mathbf{A}},{\mathbf{b}},{\mathbf{x}}_{0})$ denotes $t$ updates from lines 6-10 of Algorithm 1 for solving a general linear system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ (with exact projections).
Given matrix ${\mathbf{A}}\in\mathcal{S}_{n}^{+}$, let ${\mathbf{\Phi}}\in{\mathds{R}}^{n\times n}$ be such that ${\mathbf{A}}={\mathbf{\Phi}}{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}$. Since we denote $\bar{\mathbf{A}}={\mathbf{Q}}{\mathbf{A}}{\mathbf{Q}}^{\scriptscriptstyle{\top}}$, we can first rewrite the linear system as $\bar{{\mathbf{A}}}\bar{{\mathbf{x}}}={\mathbf{Q}}{\mathbf{b}}$ with ${\mathbf{x}}={\mathbf{Q}}^{\scriptscriptstyle{\top}}\bar{{\mathbf{x}}}$, which is further equivalent to ${\mathbf{Q}}{\mathbf{\Phi}}{\mathbf{z}}={\mathbf{Q}}{\mathbf{b}}$ with ${\mathbf{z}}={\mathbf{\Phi}}^{\scriptscriptstyle{\top}}{\mathbf{Q}}^{\scriptscriptstyle{\top}}\bar{{\mathbf{x}}}$. Thus, the preprocessing step of CD++ on ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ is equivalent to the preprocessing step of Kaczmarz++ on ${\mathbf{\Phi}}{\mathbf{z}}={\mathbf{b}}$. Denoting $\bar{{\mathbf{\Phi}}}={\mathbf{Q}}{\mathbf{\Phi}}$ and $\bar{{\mathbf{b}}}={\mathbf{Q}}{\mathbf{b}}$, let $\bar{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}{\mathbf{x}}_{t}=\mathcal{A}_{t}(\bar{{\mathbf{\Phi}}},\bar{{\mathbf{b}}},\bar{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}{\mathbf{x}}_{0})$ be the implicit Kaczmarz++ iterates, denoted as ${\mathbf{z}}_{t}=\bar{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}{\mathbf{x}}_{t}$. The projection step for these iterates is
$\bar{{\mathbf{\Phi}}}_{S}^{\scriptscriptstyle{\top}}(\bar{{\mathbf{\Phi}}}_{S}\bar{{\mathbf{\Phi}}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}(\bar{{\mathbf{\Phi}}}_{S}{\mathbf{z}}_{t}-\bar{{\mathbf{b}}}_{S})=\bar{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}{\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}(\bar{{\mathbf{A}}}_{S,S}+{\lambda}{\mathbf{I}})^{-1}(\bar{{\mathbf{A}}}_{S}{\mathbf{x}}_{t}-\bar{{\mathbf{b}}}_{S})$ (5.2)
which leads to the coordinate descent update (5.1). Notice that the iterates ${\mathbf{x}}_{t}$ of CD++ are actually solving the system $\bar{\mathbf{A}}\bar{\mathbf{x}}=\bar{\mathbf{b}}$, and thus converging to its solution, $\bar{\mathbf{x}}^{*}$, and not to ${\mathbf{x}}^{*}$. According to Corollary 4.4, letting ${\mathbf{z}}^{*}$ denote the solution of the implicit system $\bar{\mathbf{\Phi}}{\mathbf{z}}=\bar{\mathbf{b}}$, we have the following:
${\mathds{E}}\left\|{\mathbf{Q}}^{\scriptscriptstyle{\top}}{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\right\|_{{\mathbf{A}}}^{2}$ $={\mathds{E}}\left\|\bar{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}({\mathbf{x}}_{t}-\bar{\mathbf{x}}^{*})\right\|^{2}={\mathds{E}}\left\|{\mathbf{z}}_{t}-{\mathbf{z}}^{*}\right\|^{2}$
$\leq 8\Big{(}1-\frac{k}{24n\bar{\kappa}_{k}(\bar{\mathbf{\Phi}})}\Big{)}^{t}\cdot\|{\mathbf{z}}_{0}-{\mathbf{z}}^{*}\|^{2}$
$=8\Big{(}1-\frac{k}{24n\bar{\kappa}_{k}(\bar{\mathbf{\Phi}})}\Big{)}^{t}\cdot\|{\mathbf{Q}}^{\scriptscriptstyle{\top}}{\mathbf{x}}_{0}-{\mathbf{x}}^{*}\|_{{\mathbf{A}}}^{2}.$
Choosing ${\mathbf{\Phi}}={\mathbf{A}}^{1/2}$, we recover the claim. Note that the application of ${\mathbf{Q}}^{\scriptscriptstyle{\top}}$ (line 16 in Algorithm 3) transforms the iterate ${\mathbf{x}}_{t}$ back to solving the system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$. ∎
@@
Thus, the convergence rate of CD++ is determined by the convergence of the implicit Kaczmarz++ algorithm. Crucially, this convergence is now governed by
$\bar{\kappa}_{k}({\mathbf{A}}^{1/2})\leq\sqrt{\bar{\kappa}_{k}({\mathbf{A}})},$
which improves over simply running Kaczmarz++ on ${\mathbf{A}}$ by at least the square root of $\bar{\kappa}_{k}({\mathbf{A}})$. Finally, similar to Corollary 4.4, the convergence rate in Theorem 5.1 can be improved when the matrix ${\mathbf{A}}$ is rank deficient, by replacing $n$ with $2r\sqrt{\log(n\bar{\kappa}_{k})}$.
@@ 5.2 Simplified Block Memoization
This block selection strategy (which we also adapt for Kaczmarz++ in the supplement) implies that the first $B=\frac{n}{s}\log n$ blocks will be sampled uniformly at random from all size $s$ index sets and added to the block list ${\mathcal{B}}$. After that, in $T$ iterations the scheme will collect on average an additional $\sum_{t=B}^{T}\frac{B}{t}\approx B\log(\frac{T}{B})$ blocks. Since the factor $\log(\frac{T}{B})$ grows as the iterations progress, this implies that we are guaranteed to reach the number of blocks that is needed by our theory to imply convergence. However, since the factor grows slowly, we will not have to do too much unnecessary Cholesky factorizations before converging to a desired accuracy. Note that each Cholesky factor takes $O(s^{2})$ memory, so if we store $O(\frac{n}{s}\log n)$ factors throughout the convergence, then this uses only $O(ns\log n)$ additional memory.